OpenStudy (kayla6993):

Use identities to find the indicated value for each angle measure. Sin theta = 12/13, cos theta>0 find cos (2 theta).

OpenStudy (anonymous):

\[\cos(2x)=1-2\sin^2(x)\]

OpenStudy (anonymous):

in your case \[1-2\left(\frac{12}{13}\right)^2\] whatever that is